I really need your help in solving this problem. I appreciate your time and efforts. Have a terrific Evening :) Here is the question? Boxing The diagonal distance from corner to corner of a square boxing ring is 24 feet. Find the length of a side of the ring?
Answer by TimothyLamb(4379)

a = length of any side of the square
d = length of diagonal of the square
---
d = sqrt( aa + aa )
d = sqrt( 2aa )
d = sqrt( 2 ) * sqrt( aa )
d = sqrt( 2 ) * a
a = d / sqrt( 2 )
---
answer:
a = 24 / sqrt( 2 )
a = 16.97
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php